[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Aug 18, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 19 02:35:50 UTC 2017


HSU Shift Log for Fri Aug 18 2017 22:35:44 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES-16 meso sector that was supposed to begin at 00Z  
appears to have been bumped by other requests in the upper Midwest and  
northern Rockies.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170819/792b4a21/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list